Note: The old exams that follow this page are based on the first edition of the textbook, and some of the procedure names are different from the new edition. Examples:

old                new
mapcar             map
map                stream-map
head               stream-car
tail               stream-cdr
empty-stream?      stream-null?
sequence           begin
attach-type        attach-tag
type               type-tag

There may be additional minor differences; this isn’t an exhaustive list. Use the old exams as examples of the kinds of questions we ask, and the kinds of answers we want, but don’t copy code from these solutions without thinking about it!